Mr. Collins is a 55-year-old physician’s assistant with a history of chronic hepatitis C. Past history was remarkable for a needle stick injury on the job that occurred two years ago. Following the injury, he initially experienced sporadic episodes of mild flu-like symptoms with fever, chills, and diaphoresis. Over the past six months, his complaints increased to include generalized muscle weakness, fatigue, and weight loss. Mr. Collins is 5’9” tall and weighs 150 pounds with a usual weight of 162. On examination he was noted to have hepatomegaly and be markedly pale with icteric (jaundiced) conjunctiva and skin.

Mr. Collins complains of a poor appetite. He generally eats two meals a day and prefers meat, potatoes, salted snacks, and an occasional beer. He dislikes sweets, and some vegetables and fruits. He states that he often has nausea and feels bloated after eating, but is unable to pass the gas. He feels most comfortable when his stomach is empty.

1. What are the most common etiologies of hepatitis C virus (HCV)?

2. What general dietary recommendations should be given to Mr. Collins?

3. Suppose Mr. Collins develops steatorrhea. Recommend daily energy ( kcals/day)(based on his EER; light activity), protein(g/kg), and total fat(% of kcals from fat) intakes for him. show calculations, get totals for this specfic patient.
Answer #1

1. Common causes of HCV ( hepatitis C ) are :-

- child born to hepatitis C positive mother

- blood transfusion

- recieving contaminated blood products

- unsafe injection

- inadequate sterilization of medical equipment

- sexual transmission

- getting pierced with contaminated instruments

2. Diet to be given to hepatitis C positive patient are :-

Foods to be given :

- complete grains

- bread

- legumes

- beans

- vegetables

- fresh fruits

- drink lots of water

# Foods to be avoided are :

- alcohol

- sugary foods

- salty foods

- artificial sweeteners

- supplements ( only prescribed supplements to be taken )

3. Kcals /day :

Step 1: Find your body weight in kilograms (if you live in the US, just divide your weight in pounds by 2.2 to get your weight in kilograms).

= 150×0.45= 67.5kg

Step 2: Multiply your weight in kilograms by 0.9 if you are a woman or 1.0 if you are a man.

= 67.5kg × 1 = 67.5kg

Step 3: Multiply by 24

= 67.5× 24 = 1620

Step 4: Multiply by your “Lean Factor”

= 1620 × 0.85 ( lean factor for age above 28 years )

= 1377

Step 5: Multiply by Activity Modifier

= 1377× 1.55 ( Activity modifier for light activity )

= 2134.3 Kcals / day

# Protein intake for him = 0.8 gm ( minimum protein needed to stay healthy) × 67.5kg

= 54gm of protein is needed by him per day